OpenStudy (anonymous):

Ma. Matsupelleta borrowed $12.800 at an APR of 8.00% for a car. she will make 36 monthly payments how much are her monthly payments?

OpenStudy (anonymous):

First we have to find total amount for 36 months or 3 years Amount = principle (1 + rate)^ no of years i.e. A = p (1 + r)^n A = 12800 (1 + 8%)^3 A = 12800 (1 + .08)^3 A = 12800 (1.08)^3 A = 12800 (1.259712) A = 16,124.3136 divide said value by 36 you will get payment per month i.e. 16,124.3136/36 = $447.8976 for more help www.teachon.tk

OpenStudy (anonymous):

Monthly pymt = $401.11